Consider the linear system:

x + 2y + 3z = a
x + 3y + 8z = b
x + 2y + 2z = c

where a, b and c are arbitrary constants. Find all solutions of this system.

(12)

Please assist.
1
Expert's answer
2020-04-20T15:33:40-0400

From the first and the third equations it follows that "z=a-c". Substituting to the second and the third equations we have

"x+3y=b-8a+8c" and "x+2y=c-2a+2c=3c-2a".

Hence "y=b-8a+8c-3c+2a=b-6a+5c" and "x=3c-2a-2b+12a-10c=10a-2b-7c"

The answer is

"x=10a-2b-7c""y=-6a+b+5c""z=a-c"